Problem: A marketing researcher randomly selects 20 female students in a certain university and found out that their


average monthly expenditures for their cell phone loads is Php 500.00 with a standard deviation of Php75.00



a. Find the 99% confidence interval for the average expenditures of female students for their cell phone loads.



b. Find the 90% confidence interval for the average expenditures of female students for their cell phone loads.



c. Find the 80% confidence interval for the average expenditures of female students for their cell phone loads.

1
Expert's answer
2022-03-28T16:11:21-0400

Confidence interval can be found the following way:

"x \in (a-Cr_{\alpha}*{\frac {\sigma} {\sqrt n}};a+Cr_{\alpha}*{\frac {\sigma} {\sqrt n}})" , where a-sample mean, "Cr_{\alpha}" - critical value on "{\alpha}" confidence level, "\sigma" - standard deviation, n - sample size

Since the sample size is small(<30), then it is appropriate to use t-score as critical value. So, "P(T(n-1)>Cr)={\frac {1+\alpha} 2}" , where T(n-1)-Student's distribution with n-1 degrees of freedom

Which means, the sought confidence interval is "(500-Cr_{\alpha}*{\frac {75} {\sqrt {20}}};a+Cr_{\alpha}*{\frac {75} {\sqrt {20}}})=(500-16.77*Cr_{\alpha},500+16.77*Cr_{\alpha})"

a) "P(T(19)>Cr)={\frac {1+0.99} 2}=0.995\implies Cr=2.861"

So, the sought confidence interval is "(500-16.77*2.861,500+16.77*2.861)=(452, 548)"


b) "P(T(19)>Cr)={\frac {1+0.9} 2}=0.95\implies Cr=1.729"

So, the sought confidence interval is "(500-16.77*1.729,500+16.77*1.729)=(471, 529)"


c) "P(T(19)>Cr)={\frac {1+0.8} 2}=0.9\implies Cr=1.328"

So, the sought confidence interval is "(500-16.77*1.328,500+16.77*1.328)=(478, 522)"